Two particles come towards each other with speed 0.9c with respect to laboratory. Calculate their relative speeds.

1
Expert's answer
2022-03-20T18:48:55-0400

The relative velocity

"v_{r}=\frac{v+u}{1+\frac{vu}{c^2}}"

"v_r=\frac{0.9c+0.9c}{1+\frac{(0.9c)^2}{c^2}}=0.994c"